A train has a maximum speed 20m/s which it can achieve at an acceleration of 0.25m/s with its brake fully applied, the train has a deceleration of 0.5m/s². What is the shortest time the train can travel between two stations 800m apart if it's stops at both stations?.

1
Expert's answer
2022-03-31T17:23:07-0400

"t=\frac{v}{2a_1}+\frac v{2a_2}+\frac sv=100~s."